OpenStudy (kainui):

Prove or disprove for all integers n greater than 0: \(133^n\) has the same last three digits as \(375+(-1)^n 250 + 8^n\)

OpenStudy (btaylor):

Basically, we want to find \[133^n \mod 1000 = \left\{ \begin{array}{lr} 125 + (8 \mod 1000) && : n \text{ is odd}\\ 625 + (8 \mod 1000) && : n \text{ is even} \end{array} \right.\]

OpenStudy (btaylor):

Not "Basically." Definitely not basic. But that is the equation form for this problem.

OpenStudy (dan815):

from arraging a little (5^3+8)^n=133^n 375+(-1)^n*250+8^n = 3*5^3+(-1)^n*2*5^3+8^n so when n is even then 5^4+8^n when n is odd 5^3+8^n

OpenStudy (dan815):

(5^3+2^3)^n=133^n

OpenStudy (dan815):

(5^3+2^3)^n mod 10^3 (a^3+b^3)^n mod (ab)^3

OpenStudy (rational):

\[375+(-1)^n 250 + 8^n -133^n \equiv 0+(-1)^n0+8^n-8^n\equiv 0 \pmod{5^3} \] \[375+(-1)^n 250 + 8^n -133^n \equiv -1+(-1)^n2+0-5^n\equiv 0 \pmod{2^3} \] so \(2^35^3\mid 375+(-1)^n 250 + 8^n -133^n\)

OpenStudy (ikram002p):

hmm @rational how u managed 2^3 ?

OpenStudy (dan815):

ya i already diiddd it up theere :)

OpenStudy (dan815):

soo on even its 5^4 + 8^n and odd 5^3+8^n

OpenStudy (dan815):

5^4=375+250 5^3=375-250

OpenStudy (dan815):

(a^3+b^3)^n mod (ab)^3 = a^3n + b^3n mod (ab)^3

OpenStudy (ikram002p):

was typing then suddenly random music appears so i wanna type ans share\ https://www.youtube.com/watch?v=28sdV_DXSrU

OpenStudy (dan815):